In the Keynesian-cross analysis, if the consumption function is given by C = 20 + 0.7 (Y – T), and planned investment is 100, G is 100, and T is 100, then equilibrium Y is:

A) 350. B) 600. C) 950. D) 1,100.

When I calculated it, I got Y = 500. But that is not a option, Can you please walk me through the steps?
Answer #1

Ans) Aggregate demand is given by ÷

Z = C + I + G

Z = 20 + 0.7(Y-T) + I + G

At equilibrium, aggregate demand = output.

Z = Y

Y = 20 + 0.7 (Y - 100) + 100 + 100

Y = 20 + 0.7Y + 0.7(-100) + 200

Y = 220 + 0.7Y - 70

Y = 150 + 0.7 Y

Y - 0.7Y = 150

0.3Y = 150

Y = 150 ÷ 0.3

Y = 500.

(Options are not correct)